The length of a rectangle is increased by 10% and the width is decreased by 10%. find the percentage change in area.
Answer by checkley77(12569)

Original size=x^2
Modified size=1.1x*.9x=.99x^2
Difference=x^2-.99x^2=.01x^2 or a 1% decrease.